# Advancing Spatio-Temporal Processing in Spiking Neural Networks through Adaptation

### v1

@@ Abstract
Efficient implementations of spiking neural networks on neuromorphic hardware promise orders of magnitude less power consumption than their non-spiking counterparts. The standard neuron model for spike-based computation on such neuromorphic systems has long been the leaky integrate-and-fire (LIF) neuron. As a promising advancement, a computationally light augmentation of the LIF neuron model with an adaptation mechanism experienced a recent upswing in popularity, caused by demonstrations of its superior performance on spatio-temporal processing tasks. The root of the superiority of these so-called adaptive LIF neurons however, is not well understood. In this article, we thoroughly analyze the dynamical, computational, and learning properties of adaptive LIF neurons and networks thereof. We find that the frequently observed stability problems during training of such networks can be overcome by applying an alternative discretization method that results in provably better stability properties than the commonly used Euler-Forward method. With this discretization, we achieved a new state-of-the-art performance on common event-based benchmark datasets. We also show that the superiority of networks of adaptive LIF neurons extends to the prediction and generation of complex time series. Our further analysis of the computational properties of networks of adaptive LIF neurons shows that they are particularly well suited to exploit the spatio-temporal structure of input sequences. Furthermore, these networks are surprisingly robust to shifts of the mean input strength and input spike rate, even when these shifts were not observed during training. As a consequence, high-performance networks can be obtained without any normalization techniques such as batch normalization or batch-normalization through time.
@@ 1 Introduction
Although the converging evidence suggests that networks of adaptive LIF neurons are superior to LIF networks for neuromorphic applications, there are still many questions open. First, to achieve top performance, usually all neuron parameters are trained together with the synaptic weights. Changes of the neuron parameters however can quickly lead to unstable models which disrupts training. To avoid instabilities, parameter bounds have to be defined and fine-tuned. If these bounds are too wide, the network can become unstable, if they are too narrow, one cannot utilize the full computational expressivity. We show that this problem is not inherent to the neuron model but rather caused by the standard discrete-time formulation of the continuous neuron dynamics which is based on the Euler-Forward discretization method. Our thorough theoretical analysis reveals that an alternative discretization method, the so-called Symplectic-Euler discretization, provably leads to a much more stable discrete adaptive neuron model, where the full expressivity of adaptive neurons can be utilized at the same computational costs. Using this insight, we demonstrate the power of adaptation by showing that our improved adaptive RSNNs outperform the state-of-the-art on spiking speech recognition datasets as well as an ECG dataset. We then show that the superiority of adaptive RSNNs is not limited to classification tasks but extends to the prediction and generation of complex time series.
Second, there is a lack of understanding why sub-threshold adaptation is so powerful in RSNNs. We thoroughly analyze the computational dynamics in single adaptive LIF neurons, as well as in networks of such neurons. Our analysis suggests that adaptive LIF neurons are especially capable of detecting temporal changes in input spike density, while being robust to shifts of total spike rate. Hence, adaptive RSNNs are well suited to analyze the temporal properties of input sequences.
Third, high-performance SNNs are usually trained using normalization techniques such as batch normalization or batch normalization through time [batchnorm, kim2021revisitingbatchnormalizationtraining, deckers2024co, vicente-solaSpikingNeuralNetworks2023]. These methods however complicate the training process and the implementation of networks on neuromorphic hardware. We show that adaptation has a previously unrecognized benefit on network optimization. Since adaptation inherently stabilizes network activity, we hypothesized that explicit normalization is not necessary in adaptive RSNNs. In fact, all our results were obtained without explicit normalization techniques. We test this hypothesis and show that in contrast to LIF networks, networks of adaptive LIF neurons can tolerate substantial shifts in the mean input strength as well as substantial levels of background noise even when these perturbations were not observed during training.
2 Results
2.1 Adaptive LIF neurons
@@
Figure 2: The adLIF neuron model shows frequency-selective resonating behavior. a) Voltage response (root mean squared membrane potential over $10$ seconds) of a LIF neuron in response to tonic spike input of different rates. b) Membrane potential response of an adLIF neuron with intrinsic frequency of $f\approx 60$ Hz to an input spike triplet at $60$ Hz (left) and $100$ Hz (right). c) Voltage response of the same adLIF neuron as in panels b and c for tonic spiking input at different rates. d) A $10$ Hz sinusoidal input signal (top) is encoded as an input spike train (middle) through spike frequency modulation (SFM), see main text for details. Membrane potential response of an adLIF neuron with intrinsic frequency $f\approx 10$ Hz (bottom). e) Same as panel e, but for a sinusoidal input at $7$ Hz. f) Voltage response of an adLIF neuron to SFM-encoded sinusoidal input at various frequencies. g) Same as panel f, but for a LIF neuron. See Methods for parameters and input generation.
@@
We took this analysis a step further and asked whether this resonance could account for frequencies in the input spike train that are not directly encoded by spike rate, but rather by slow changes of the spike rate over time, a coding scheme previously termed spike frequency modulation (SFM) sfmwang. As a guiding example, we encoded a slow-varying sinusoidal signal as a spike train, where the magnitude of the signal at a certain time is given by the local spike rate, shown in Fig. 2d. The spike rate thereby varied between 00 Hz and $200$ Hz, whereas the underlying, encoded sinus signal oscillated with a constant frequency of $10$ Hz. Again, we see increased response of the membrane potential $u$ over time in the case of the $10$ Hz input compared to a slower $7$ Hz sinus signal (Fig. 2d,e), due to resonance with the adLIF neuron, see also Fig. 2f. In contrast, the corresponding membrane voltage response amplitude of a LIF neuron is almost indifferent to the intrinsic frequency of the underlying sinus input, see Fig. 2g. This shows that in contrast to the LIF neuron, the adLIF neuron model is sensitive to the longer-term temporal structure, i.e. variation of the input signal. In Section Computational properties of adLIF networks, we highlight the importance of this frequency-dependence of neuron responses as a key ingredient for the powerful feature detection capabilities of networks of adLIF neurons.
2.2 The Symplectic-Euler discretized adLIF neuron
@@
$w[k]$ $=\beta w[k-1]+(1-\beta)(au[k-1]+bS[k]),$ (6b)
@@
We refer to this neuron model as the SE-adLIF model in order to distinguish it from the Euler-Forward discretized model. Note that the reset mechanism from Eq. (8) is applied to obtain $u[k]$ from $\hat{u}[k]$ before computing $w[k]$. While it is also possible to apply the reset after computing $w[k]$, we found that the above described way yields the best performance. For the sub-threshold dynamics, this leads to update matrices (see Section 4.2 in Methods)
@@
2.3 Stability analysis of discretized adLIF models
@@
Training of networks of adLIF neurons often suffers from stability issues, in particular when the neuron parameters such as the time constants $\tau_{w}$, $\tau_{u}$ or the adaptation parameters $a$, $b$ are trained. The decay rate $r$ (see Fig. 1b) determines the asymptotic stability of the neuron dynamics: A decay rate $r>1$ results in a neuron in which the state vector grows exponentially over time, leading to an unstable and diverging system, whereas a neuron with decay rate $r<1$ exhibits stable dynamics where the state vector exponentially decays to a resting state equilibrium. Since the continuous-time adLIF neuron model Eq. (2) is inherently stable for $a\geq-1$ (see Section 4.3 in Methods for proof), it is a desirable property of a discretization method to preserve this stability for all possible parameterizations. However, when the model is time-discretized, the discrete-time decay rate can deviate from its continuous counterpart due to discretization errors, leading to stability properties that deviate from the continuous-time system. This is shown empirically in Fig. 3a. We instantiated $1{,}000$ different neurons for both discretization methods, Euler-Forward and SE, as well as the continuous model, in a grid-like manner over a reasonable parameter range and plot their calculated frequencies and decay rates in Fig. 3a. While the continuous system is stable for all considered parameter combinations (middle panel), the Euler-Forward approximation (left panel) is unstable for many parameterizations (decay rate $r>1$). In contrast, for the SE discretization (right panel), all parameter combinations resulted in stable neuron dynamics ($r<1$). These empirical results show that the SE discretization more closely follows the stability properties of the continuous model, whereas the Euler-Forward method deviates drastically from both. How can this discrepancy be explained?
@@
When analyzing the discretized neurons, one has to calculate $f$ and $r$ directly from the discrete system by calculating the eigenvalues $\lambda_{1,2}$ of the state transition matrix $\bar{A}$. This allows to study the behavior and the stability of the adLIF neuron model for different discretizations. Two cases have to be differentiated: If the eigenvalues are complex, the membrane potential exhibits oscillations, whereas if they are real, no oscillations occur and the neuron behaves similar to a LIF neuron. In the complex case, we can write the eigenvalues in polar form as $\lambda_{1,2}=re^{\pm j\phi}$, where $j$ denotes the imaginary unit. Hence, the eigenvalues are complex conjugates, the decay rate $r$ is given by their modulus (see Fig. 3b) and $\phi$ is obtained as the argument of $\lambda_{1}$ ($\text{arg}(\lambda_{2})=-\phi$). Intuitively, the angle $\phi$ is the rotation of the neuron state with each time step $\Delta t$ in radians, and hence determines the frequency of the oscillation. One thus obtains the intrinsic frequency $f$ in Hertz as $f=\frac{\phi}{2\pi\Delta t}$. In the case of real eigenvalues, $r$ is given by the magnitude of the largest eigenvalue. AdLIF neurons can thereby represent underdamped (complex eigenvalues), critically damped (equal real eigenvalues), and overdamped (non-equal real eigenvalues) systems via different parameterizations. Note, that only in the underdamped case the neuron can oscillate.
@@
Figure 3: Stability of adLIF discretizations. a) Intrinsic frequency $f$ and per-timestep decay rate $r$ of $1{,}000$ different parameterizations of adLIF neurons for Euler-Forward discretization (left), SE (right), and the continuous model (middle). The horizontal dotted line at $r=1.0$ marks the stability bound. Instances above this line diverge due to exponential growth. Parameter ranges are uniformly distributed over the intervals $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. b) Eigenvalues of $\bar{A}_{\mathrm{EF}}$ (left) and $\bar{A}_{\mathrm{SE}}$ (right) plotted in the complex plane for fixed $\tau_{u}=25$ ms, $\tau_{w}=60$ ms and varying $a\in[10,800]$. Decay rate $r$ as modulus of the eigenvalue $\lambda_{1}$ and angle $\phi$ as argument of $\lambda_{1}$ are shown for $a=282$, marked with * and ** for EF and SE respectively. The grey half-circle denotes the stable region of $r\leq 1$. c) Relationship of parameter $a$ to intrinsic frequency $f$ (top) and decay rate $r$ (bottom) for the same $\tau_{u}$ and $\tau_{w}$ as in panel b. Points with * and ** denote the corresponding eigenvalues from panel b. Recall the linear relationship $f=\frac{\phi}{2\pi\Delta t}$ between angle $\phi$ and $f$ of the discrete models. Horizontal grey line in bottom panel denotes stability boundary of $r=1$. Values for $r$ of continuous model ($r=0.897$) and SE-adLIF ($r=0.887$) are constant w.r.t. $a$ (SE-adLIF values for $r$ not visible due to near-perfect fit to the continuous model). d) Maximum admissible frequency for stable dynamics for Euler-Forward discretization with $\Delta t=1$ ms over different values of $\tau_{u}$ and $\tau_{w}$, where $a$ is set to the maximum stable value $a_{\mathrm{max}}$ (see main text and Section 4.7 in Methods).
@@
As described above, the eigenvalues $\lambda_{1,2}$ of state transition matrices $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ are determining the stability of the discrete neurons. We can directly observe the origin of instability for the EF-adLIF by plotting eigenvalues for different neuron parameters in the complex plane. In Fig. 3b, we show some eigenvalues of $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ in the case of fixed time constants $\tau_{u}$ and $\tau_{w}$ and varying parameter $a$. In the complex plane, the stability boundary appears as circle, separating the stable ($r<1$, grey area) from the unstable region ($r>1$). Our analysis in Section 4.4 in Methods shows that for the Euler-Forward method, for fixed time constants $\tau_{u}$ and $\tau_{w}$, the real part $\Re(\lambda_{1,2})$ of these eigenvalues is constant and strictly positive with respect to $a$, such that the eigenvalues are aligned along a vertical line in the right half-plane of the complex plane (see Fig. 3b, left panel). As $a$ increases, the imaginary part increases and so does the decay rate $r$. Hence, for the Euler-Forward discretized neuron, the eigenvalues overshoot the stability boundary $r=1$ for increasing $a$. This leads to a drastically reduced range of the angle $\phi$ and therefore a reduced range of the intrinsic frequency $f$ where the neuron is stable. In contrast, for the SE discretized model, the parameter $a$ controls only the angle $\phi$ of the eigenvalues (see Section 4.5 in Methods) and hence the intrinsic frequency $f$, but not the decay rate $r$. The decay rate is given by $r=\sqrt{\alpha\beta}$ (see Eq. (55)) and is hence guaranteed to stay within the stability bound $r<1$ for all $\tau_{w}>0$ and $\tau_{u}>0$, see Fig. 3b, right panel and Fig. 3c.
@@
We analytically calculated stability bounds for both the Euler-Forward and SE discretization, see Methods Sections 4.4 and 4.5. For each given tuple of time constants $\tau_{u}$ and $\tau_{w}$ we calculated a corresponding $a_{\mathrm{max}}$, that is, the maximum value of the parameter $a$ for which the model is still stable. This analysis shows that the SE discretization allows the neuron to utilize the full frequency bandwidth up to the Nyquist frequency at $\frac{1}{2\Delta t}$, at which aliasing occurs. Since we used a discretization time step of $\Delta t=1$ ms for Fig. 3, the Nyquist frequency is $500$ Hz. Theorem 2.1 below summarizes the full frequency coverage of SE-adLIF and the stability within this frequency range (see Methods, Section 4.6 for a proof).

### v3

@@ Abstract
Implementations of spiking neural networks on neuromorphic hardware promise orders of magnitude less power consumption than their non-spiking counterparts. The standard neuron model for spike-based computation on such systems has long been the leaky integrate-and-fire (LIF) neuron. A computationally light augmentation of the LIF neuron model with an adaptation mechanism has recently been shown to exhibit superior performance on spatio-temporal processing tasks. The root of the superiority of these so-called adaptive LIF neurons however is not well understood. In this article, we thoroughly analyze the dynamical, computational, and learning properties of adaptive LIF neurons and networks thereof. Our investigation reveals significant challenges related to stability and parameterization when employing the conventional Euler-Forward discretization for this class of models. We report a rigorous theoretical and empirical demonstration that these challenges can be effectively addressed by adopting an alternative discretization approach – the Symplectic Euler method, allowing to improve over state-of-the-art performances on common event-based benchmark datasets. Our further analysis of the computational properties of networks of adaptive LIF neurons shows that they are particularly well suited to exploit the spatio-temporal structure of input sequences without any normalization techniques.
@@ 1 Introduction
Although the converging evidence suggests that networks of adaptive LIF neurons are superior to LIF networks for neuromorphic applications, there are still many questions open. First, to achieve top performance, usually all neuron parameters are trained together with the synaptic weights. Changes of the neuron parameters however can quickly lead to unstable models which disrupts training. To avoid instabilities, parameter bounds have to be defined and fine-tuned. If these bounds are too wide, the network can become unstable, if they are too narrow, one cannot utilize the full computational expressivity.
In this work, we show that this problem is not inherent to the neuron model but rather caused by the standard discrete-time formulation of the continuous neuron dynamics which is based on the Euler-Forward discretization method. Despite mere stability issues, we identified a plethora of drawbacks arising from the application of the widely used Euler-Forward discretization. These include, inter alia, unintended interdependencies between neuron parameters, deviations of the discrete model dynamics from its continuous counterpart, limitations in neuron expressibility, strong dependence between the discretization time step and the neuron dynamics, non-trivial divergence boundaries. Our thorough theoretical analysis reveals that the alternative, evenly simple Symplectic-Euler discretization method remarkably alleviates the drawbacks of the Euler-Forward method almost entirely, without additional computational cost or implementation complexity. While this discretization method could in principle be applied to an entire family of multi-dimensional neuron models, we mainly focus our theoretical and empirical analyses on a specific adaptive neuron model that recently gained traction. Using this insight, we demonstrate the power of adaptation by showing that our improved adaptive RSNNs outperform the state-of-the-art on spiking speech recognition datasets as well as an ECG dataset. We then show that the superiority of adaptive RSNNs is not limited to classification tasks but extends to the prediction and generation of complex time series. Second, there is a lack of understanding why sub-threshold adaptation is so powerful in RSNNs. We thoroughly analyze the computational dynamics in single adaptive LIF neurons, as well as in networks of such neurons. Our analysis suggests that adaptive LIF neurons are especially capable of detecting temporal changes in input spike density, while being robust to shifts of total spike rate. Hence, adaptive RSNNs are well suited to analyze the temporal properties of input sequences. Third, high-performance SNNs are usually trained using normalization techniques such as batch normalization or batch normalization through time [guoMembranePotentialBatch2023, jiangTABTemporalAccumulated2023, zhengGoingDeeperDirectlyTrained2021, vicente-solaSpikingNeuralNetworks2023]. These methods however complicate the training process and the implementation of networks on neuromorphic hardware. We show that adaptation has a previously unrecognized benefit on network optimization. Since adaptation inherently stabilizes network activity, we hypothesized that explicit normalization is not necessary in adaptive RSNNs. In fact, all our results were obtained without explicit normalization techniques. We test this hypothesis and show that in contrast to LIF networks, networks of adaptive LIF neurons can tolerate substantial shifts in the mean input strength as well as substantial levels of background noise even when these perturbations were not observed during training.
Results
Adaptive LIF neurons
@@ Results
Figure 2: The adLIF neuron model shows frequency-selective resonating behavior. a) Voltage response (root mean squared membrane potential over $10$ seconds) of a LIF neuron in response to tonic spike input of different rates. b) Membrane potential response of an adLIF neuron with intrinsic frequency of $f\approx 60$ Hz to an input spike triplet at $60$ Hz (left) and $100$ Hz (right). c) Voltage response of the same adLIF neuron as in panels b and c for tonic spiking input at different rates. d) A $10$ Hz sinusoidal input signal (top) is encoded as an input spike train (middle) through spike frequency modulation (SFM), see main text for details. Membrane potential response of an adLIF neuron with intrinsic frequency $f\approx 10$ Hz (bottom). e) Same as panel e, but for a sinusoidal input at $7$ Hz. f) Voltage response of an adLIF neuron to SFM-encoded sinusoidal input at various frequencies. g) Same as panel f, but for a LIF neuron. See ”Methods” for parameters and input generation.
@@
We took this analysis a step further and asked whether this resonance could account for frequencies in the input spike train that are not directly encoded by spike rate, but rather by slow changes of the spike rate over time, a coding scheme previously termed spike frequency modulation (SFM) sfmwang. As a guiding example, we encoded a slow-varying sinusoidal signal as a spike train, where the magnitude of the signal at a certain time is given by the local spike rate, shown in Fig. 2d. The spike rate thereby varied between 00 Hz and $200$ Hz, whereas the underlying, encoded sinus signal oscillated with a constant frequency of $10$ Hz. Again, we see increased response of the membrane potential $u$ over time in the case of the $10$ Hz input compared to a slower $7$ Hz sinus signal (Fig. 2d,e), due to resonance with the adLIF neuron, see also Fig. 2f. In contrast, the corresponding membrane voltage response amplitude of a LIF neuron is almost indifferent to the intrinsic frequency of the underlying sinus input, see Fig. 2g. This shows that in contrast to the LIF neuron, the adLIF neuron model is sensitive to the longer-term temporal structure, i.e. variation of the input signal. In Section ”Computational properties of adLIF networks”, we highlight the importance of this frequency-dependence of neuron responses as a key ingredient for the powerful feature detection capabilities of networks of adLIF neurons.
The Symplectic-Euler discretized adLIF neuron
@@
$w[k]$ $=\beta w[k-1]+(1-\beta)\left(au[k-1]+bS[k]\right),$ (6b)
@@
We refer to this neuron model as the SE-adLIF model in order to distinguish it from the Euler-Forward discretized model. Note that the reset mechanism from Eq. (8) is applied to obtain $u[k]$ from $\hat{u}[k]$ before computing $w[k]$. While it is also possible to apply the reset after computing $w[k]$, we found that the above described way yields the best performance. For the sub-threshold dynamics, this leads to update matrices (see Section ”Derivation of matrices $\bar{A}_{\mathrm{SE}}$ and $\bar{B}_{\mathrm{SE}}$ for the SE-adLIF neuron” in ”Methods”)
@@
Stability analysis of discretized adLIF models
@@
A desirable characteristic of a discretization method is its ability to maintain a close alignment between the discretized system and the continuous ground-truth. In contrast to SE, the EF discretization exhibits a pronounced dependence of this alignment on the discretization time step $\Delta t$, thereby reducing its robustness. This is visualized in Fig. 3a, where we discretized an adLIF neuron with the EF method (left) and the SE method (right) using $3$ different discretization time steps $\Delta t$. We observed that the EF-discretized neuron clearly diverges for larger values of $\Delta t$, in this example even for $\Delta t=1$ ms. In contrast, the same neuron discretized with the Symplectic Euler method is robust to the choice of $\Delta t$. The divergence of the neuron can be quantified by its decay rate $r$, which gives the exponential decay of the envelope of a neuron’s membrane potential $u(t)$ (see also Fig. 1b). For $r<1$, the neuron stably decays to a resting-state equilibrium. However, if this decay rate exceeds $1$, the neuron becomes unstable and its membrane potential grows indefinitely, as observable for the EF-adLIF neuron with $\Delta t=1$ in Fig. 3a. When comparing the relationship of the decay rate $r$ with respect to discretization time step $\Delta t$, as visualized in Fig. 3b, the favorable adherence of the SE-discretized adLIF to the continuous model is evident. The SE-adLIF decay rate is independent of $\Delta t$ and evaluates to $r\approx 0.972$, which is the decay rate of the continuous model. For the EF discretization in contrast, $r$ grows along with increasing $\Delta t$, resulting in discretized neurons exceeding the stability boundary at $r=1$. As the computational cost of training SNNs via BPTT increases with smaller discretization time steps due to longer sequence lengths, the SE discretization is clearly favorable over EF, since it ensures stability and adherence to the continuous ground truth when $\Delta t$ is large. The adherence of SE-adLIF to the continuous adLIF model is not limited to its robustness to the choice of $\Delta t$. For a given $\Delta t=1$ ms, SE-adLIF follows the characteristics of the continuous model with respect to its parameters $\tau_{u}$, $\tau_{w}$, and $a$ much closer. While time constants $\tau_{u}$ and $\tau_{w}$ affect the decay rate $r$ of the adLIF neuron, parameter $a$ determines the frequency of oscillation. Since the continuous-time adLIF neuron model Eq. (2) is inherently stable for $a\geq-1$ (see Section ”Proof of stability bounds for the continuous adLIF model” in ”Methods” for proof), it is a desirable property of a discretization method to preserve this stability for all possible parameterizations. This is shown empirically in Fig. 3c. We instantiated $1{,}000$ different neurons for both discretization methods, Euler-Forward and SE, as well as the continuous model, in a grid-like manner over a reasonable parameter range and plot their calculated frequencies and decay rates in Fig. 3c. While the continuous system is stable for all considered parameter combinations (middle panel), the Euler-Forward approximation (left panel) is unstable for many parameterizations (decay rate $r>1$). In contrast, for the SE discretization (right panel), all parameter combinations resulted in stable neuron dynamics ($r<1$). These empirical results show that the SE discretization more closely follows the stability properties of the continuous model, whereas the Euler-Forward method deviates drastically from both. How can this discrepancy be explained?
@@
When analyzing the discretized neurons, one has to calculate $f$ and $r$ directly from the discrete system by calculating the eigenvalues $\lambda_{1,2}$ of the state transition matrix $\bar{A}$. This allows to study the behavior and the stability of the adLIF neuron model for different discretizations. Two cases have to be differentiated: If the eigenvalues are complex, the membrane potential exhibits oscillations, whereas if they are real, no oscillations occur and the neuron behaves similar to a LIF neuron. In the complex case, we can write the eigenvalues in polar form as $\lambda_{1,2}=re^{\pm j\phi}$, where $j$ denotes the imaginary unit. Hence, the eigenvalues are complex conjugates, the decay rate $r$ is given by their modulus (see Fig. 3d) and $\phi$ is obtained as the argument of $\lambda_{1}$ ($\text{arg}(\lambda_{2})=-\phi$). Intuitively, the angle $\phi$ is the rotation of the neuron state with each time step $\Delta t$ in radians, and hence determines the frequency of the oscillation. One thus obtains the intrinsic frequency $f$ in Hertz as $f=\frac{\phi}{2\pi\Delta t}$. In the case of real eigenvalues, $r$ is given by the magnitude of the largest eigenvalue. AdLIF neurons can thereby represent underdamped (complex eigenvalues), critically damped (equal real eigenvalues), and overdamped (non-equal real eigenvalues) systems via different parameterizations. Note, that only in the underdamped case the neuron can oscillate.
@@
Figure 3: Stability of adLIF discretizations. a) Membrane potential $u(t)$ over time for an EF-adLIF (left) and a SE-adLIF (right) neuron for different discretization time steps $\Delta t\in\{0.001,0.5,1\}$. Both neurons have the same parameters ($\tau_{u}=25$ ms, $\tau_{w}=60$ ms, $a=120$). b) Relationship between the decay rate $r$ and discretization time step $\Delta t$ for adLIF models with different discretizations, EF and SE. All decay rates are calculated with respect to $1$ ms, a decay rate of $r=0.9$ hence represents a decrease in magnitude of $10\%$ every $1$ ms. The decay rate ($r=0.972$) of the equivalently parameterized continuous model is highlighted. Same neuron parameters as in panel a. c) Intrinsic frequency $f$ and per-timestep decay rate $r$ of $1{,}000$ different parameterizations of adLIF neurons for Euler-Forward discretization (left), SE (right), and the continuous model (middle). The horizontal dotted line at $r=1.0$ marks the stability bound. Instances above this line diverge due to exponential growth. Parameter ranges are uniformly distributed over the intervals $a\in[0,120]$, $\tau_{u}\in[5,25]$ ms and $\tau_{w}\in[60,300]$ ms. d) Eigenvalues of $\bar{A}_{\mathrm{EF}}$ (left) and $\bar{A}_{\mathrm{SE}}$ (right) plotted in the complex plane for fixed $\tau_{u}=25$ ms, $\tau_{w}=60$ ms and varying $a\in[10,800]$. Decay rate $r$ as modulus of the eigenvalue $\lambda_{1}$ and angle $\phi$ as argument of $\lambda_{1}$ are shown for $a=282$, marked with * and ** for EF and SE respectively. The grey half-circle denotes the stable region of $r\leq 1$. e) Relationship of parameter $a$ to intrinsic frequency $f$ (top) and decay rate $r$ (bottom) for the same $\tau_{u}$ and $\tau_{w}$ as in panel d. Points with * and ** denote the corresponding eigenvalues from panel b. Recall the linear relationship $f=\frac{\phi}{2\pi\Delta t}$ between angle $\phi$ and $f$ of the discrete models. Horizontal grey line in bottom panel denotes stability boundary of $r=1$. Values for $r$ of continuous model ($r=0.897$) and SE-adLIF ($r=0.887$) are constant w.r.t. $a$ (SE-adLIF values for $r$ not visible due to near-perfect fit to the continuous model). f) Maximum admissible frequency for stable dynamics for Euler-Forward discretization with $\Delta t=1$ ms over different values of $\tau_{u}$ and $\tau_{w}$, where $a$ is set to the maximum stable value $a_{\mathrm{max}}$ (see main text and Section ”Stable ranges for intrinsic frequencies of EF-adLIF” in ”Methods”).
@@
As described above, the eigenvalues $\lambda_{1,2}$ of state transition matrices $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ are determining the stability of the discrete neurons. We can directly observe the origin of instability for the EF-adLIF by plotting eigenvalues for different neuron parameters in the complex plane. In Fig. 3d, we show some eigenvalues of $\bar{A}_{\mathrm{EF}}$ and $\bar{A}_{\mathrm{SE}}$ in the case of fixed time constants $\tau_{u}$ and $\tau_{w}$ and varying parameter $a$. In the complex plane, the stability boundary appears as circle, separating the stable ($r<1$, grey area) from the unstable region ($r>1$). Our analysis in Section ”Derivation of stability bounds for EF-adLIF” in ”Methods” shows that for the Euler-Forward method, for fixed time constants $\tau_{u}$ and $\tau_{w}$, the real part $\Re(\lambda_{1,2})$ of these eigenvalues is constant and strictly positive with respect to $a$, such that the eigenvalues are aligned along a vertical line in the right half-plane of the complex plane (see Fig. 3d, left panel). As $a$ increases, the imaginary part increases and so does the decay rate $r$. As already mentioned, in the continuous adLIF model, parameter $a$ only affects the frequency of oscillation, but not the decay rate. The SE-adLIF model adheres to this property, since the modulus of the eigenvalues does not change with respect to $a$. For EF-adLIF however, parameter $a$ exhibits an undesired side-effect on the modulus. Hence, for the Euler-Forward discretized neuron, the eigenvalues overshoot the stability boundary $r=1$ for increasing $a$. This leads to a drastically reduced range of the angle $\phi$ and therefore a reduced range of the intrinsic frequency $f$ where the neuron is stable. In contrast, for the SE discretized model, the parameter $a$ controls only the angle $\phi$ of the eigenvalues (see Section ”Derivation of the stability bounds of SE-adLIF” in ”Methods”) and hence the intrinsic frequency $f$, but not the decay rate $r$. The decay rate is given by $r=\sqrt{\alpha\beta}$ (see Eq. (55)) and is hence guaranteed to stay within the stability bound $r<1$ for all $\tau_{w}>0$ and $\tau_{u}>0$, see Fig. 3d, right panel and Fig. 3e.
@@
We analytically calculated stability bounds for both the Euler-Forward and SE discretization, see ”Methods” Sections ”Derivation of stability bounds for EF-adLIF” and Derivation of the stability bounds of SE-adLIF. For each given tuple of time constants $\tau_{u}$ and $\tau_{w}$ we calculated a corresponding $a_{\mathrm{max}}$, that is, the maximum value of the parameter $a$ for which the model is still stable. This analysis shows that the SE discretization allows the neuron to utilize the full frequency bandwidth up to the Nyquist frequency at $\frac{1}{2\Delta t}$, at which aliasing occurs. Since we used a discretization time step of $\Delta t=1$ ms for Fig. 3, the Nyquist frequency is $500$ Hz. Theorem 1.1 below summarizes the full frequency coverage of SE-adLIF and the stability within this frequency range (see ”Methods”, Section ”Proof of Theorem 1.1” for a proof).